A light plane flew from its home base to an airport 255 miles away. With a headwind, the trip took 1.7 hours. The return trip with tailwind took 1.5 hours. Fine the average airspeed of the plane and the average wind speed.
Answer by checkley71(8403)

255/1.7=150 MPH FOR THE SPEED WITH A HEAD WIND.
255/1.5=170 MPH FOR THE TRIP WITH THE WIND.
THE DIFFERENCE IN SPEEDS DIVIDED BY 2 GIVES US THE WIND SPEED.
(170-150)/2=20/2=10 MPH FOR THE WIND SPEED.
170-10=160 FOR THE PLANE SPEED.
150+10=160